Number of lines in paschen series when electron return from mna nth energy level?

Grade:12th pass

1 Answers

Piyush Kumar Behera
5 years ago
@nishant
Paschen series means electron returns to 3rd orbit.
Means effectively we have to find out no.of ways of returning from nth orbit to 3rd orbit.
The no. of lines from nth orbit to the third orbit is same as the no.of orbit from n-2 orbit to 1st orbit.This can be understood by simple analysis or logical thinking.
we know the formula(I am not going to derivation,if any one wants you can write below,I will derive it) for no. of spectral lines from nth orbit to 1st orbit is
                                                =\frac{n(n-1)}{2}
as here we are going from n-2 th to 1st  orbit so the number of spectral lines is
                                        =\frac{(n-2)(n-3)}{2}
                                  
 
Hence the answer is   \frac{(n-2)(n-3)}{2}.
Hope it helps!!
Regards
Piyush Kumar Behera